OpenStudy (brightemmerald):

If (20, 14) is an ordered pair of the inverse of F(x), which of the following is an ordered pair of the function F(x)?

OpenStudy (mathmale):

All you need to do is to reverse the order of the coordinates 20 and 14.

OpenStudy (mathmate):

Hint: If g(x)=x+4, then (3,7) is an ordered pair of g(x). Find the inverse of g(x) :\(g^{-1}(x)\) 1. x=\(g^{-1}(x)\)+4 2. \(g^{-1}(x)\)=x-4 So \(g^{-1}(7)=7-4=3\) which means that (7,3) is an ordered pair of \(g^{-1}(x)\).